OpenStudy (imtant):

How to integrate y-3 / y?

sam (.sam.):

\[\int\limits \frac{y-3}{y}dy?\]

OpenStudy (imtant):

Yes.

sam (.sam.):

Separate the denominator

sam (.sam.):

\[\int\limits 1-\frac{3}{y}dy\]

sam (.sam.):

\[\int\limits 1-3y^{-1}dy\] Its just power rule now

OpenStudy (imtant):

So I wil get y-3y..right?

sam (.sam.):

Nope, integrating \(y^{-1}\) gives off ln(y)

OpenStudy (imtant):

It is y - 3 In y?

sam (.sam.):

Careful, you forgot the constant

OpenStudy (imtant):

Is it correct?

sam (.sam.):

Nope, I told you already you forgot the constant. You will get a constant whenever you do indefinite integrals

OpenStudy (imtant):

So 1-3 In y?The constant is 1,right?Is it?

sam (.sam.):

No the constant is "c" \[\int\limits\limits 1-3y^{-1}dy\] \[=y-3\ln(y)+c\]

OpenStudy (imtant):

Okay,I got it. But if the question also have dx and dy,where must I put C?